# Sparse Autoencoder Neural Operators: Model Recovery in Function Spaces

### v2

@@ 2 SAE Neural Operators for Model Recovery in Function Spaces
The goal of model recovery with SAE-NO in function spaces is to infer a sparse code $\bm{z}$} as a sample from the function space $\mathcal{Z}$, representing the data sample from the function space $\mathcal{X}$ via the operator $\mathcal{G}_{{\bm{D}}}$: $\quad\text{(encoder)}\quad{\bm{z}}=\mathcal{H}_{\theta}({\bm{x}}),\quad\text{(decoder)}\quad\hat{\bm{x}}=\mathcal{G}_{{\bm{D}}}({\bm{z}})$, where $\mathcal{H}_{\theta}$ is the encoder operator, and model recovery involves estimating the concepts $\bm{D}$}.
We consider Fourier parametrization and define SAE Fourier Neural Operators (SAE-FNOs) as $\mathcal{G}_{{\bm{D}}}({\bm{z}})\coloneqq\mathcal{F}^{-1}\big(\sum_{c=1}^{C}{\bm{W}}_{c}\cdot(\mathcal{F}{\bm{z}}_{c})\big)$, where ${\bm{W}}_{c}=\mathcal{F}{\bm{D}}_{c}$ is the Fourier transform of the convolution kernel. This reduces sparse functional model recovery to recovery of a sparse convolutional generative model (Def.˜B.1), implemented as a sum of convolutions between $\bm{z}$} and $\bm{D}$}. This formulation generalizes SAEs to operator and Fourier domains, enabling concept learning in function spaces. We further extend this approach to model recovery in lifted spaces (Def.˜2.4), incorporating lifting and projection modules that are crucial components of neural operators [12, 13, 14].
Definition 2.4 (Lifted SAE-NO for Model Recovery).
@@
The goal of model recovery with lifted SAE-NO in function spaces is to infer a sparse code $\bm{z}$} as a sample from the function space $\mathcal{Z}$, representing the data from $\mathcal{X}$ via the operator $\mathcal{G}_{{\bm{D}}_{{\bm{L}}}}$:
(lifting and encoder) ${\bm{z}}=\mathcal{H}_{\theta}({\bm{y}}),\quad{\bm{y}}={\bm{L}}{\bm{x}}$ (3)
(decoder and projecting) $\hat{{\bm{x}}}={\bm{P}}\hat{{\bm{y}}},\quad\hat{{\bm{y}}}=\mathcal{G}_{{\bm{D}}_{{\bm{L}}}}({\bm{z}}),$
@@
where $\mathcal{H}_{\theta}$ encodes the lifted data $y$, and model recovery is done via the learned lifted concepts ${\bm{D}}_{{\bm{L}}}$.
Finally, we informally define architectural inference equivalence, i.e., when two architectures can encapsulate the same underlying model and produce the same latent representations of an input.
@@ 3 Results
We generate data under the sparse convolutional generative model, and study the impact of lifting and operator on model recovery on SAE-CNN, L-SAE-CNN, SAE-FNO, and L-SAE-FNO (see Appendix). The concepts in SAE-CNN (or L-SAE-CNN), SAE-FNO (or L-SAE-FNO) are parameterized as a convolution and a Fourier-based operator, respectively.
@@
Lifting We examined model recovery in the lifted space and found: i) models can be recovered in the lifted space (Fig.˜4); ii) lifting can act as a preconditioner (Prop.˜3.1), accelerating recovery (Fig.˜4) by reducing dictionary atom correlation, promoting a more isotropic loss landscape (Fig.˜6); and iii) when lifting and projection are tied and orthogonal, the architectural inference (Props.˜3.2 and D.1) and training dynamics (Props.˜3.1 and D.2) of lifted SAEs reduces to SAEs (Fig.˜5).
Proposition 3.1 (Training Dynamics of Lifting).
The training dynamics of the lifted-SAE (L-SAE) ${\bm{D}}_{L}^{(k+1)}={\bm{D}}_{L}^{(k)}+\eta_{L}{\bm{P}}^{\top}({\bm{x}}-{\bm{P}}{\bm{D}}_{L}^{(k)}{\bm{z}}){\bm{z}}^{\top}$, with lifting $\bm{L}$} and projection $\bm{P}$}, has the effective update in the original space, expressed as: ${\bm{D}}^{(k+1)}={\bm{D}}^{(k)}+\eta_{L}({\bm{L}}^{\top}{\bm{L}})({\bm{x}}-{\bm{D}}^{(k)}{\bm{z}}){\bm{z}}^{\top}$, where ${\bm{L}}^{\top}{\bm{L}}$ acts as a preconditioner, potentially accelerating learning by inducing a more isotropic update. If they satisfy the equivalent architectural inference (Prop.˜3.2), then the dynamics are equivalent to those of an SAE: ${\bm{D}}^{(k+1)}={\bm{D}}^{(k)}+\eta({\bm{x}}-{\bm{D}}^{(k)}{\bm{z}}){\bm{z}}^{\top}$ (see Prop.˜D.2 for L-SAE-CNN).
Proposition 3.2 (Architectural Inference Equivalence of Lifting in SAE).
The architectural inference of a SAE is equivalent to a lifted-SAE if the projection is tied to the lifting operator (${\bm{P}}={\bm{L}}^{\top}$), and the lifting operator is orthogonal (${\bm{L}}^{\top}{\bm{L}}={\bm{I}}$), i.e., under the same learned model ${\bm{D}}={\bm{P}}{\bm{D}}_{L}$, the representation dynamics and output of the networks are equivalent (see Prop.˜D.1 for SAE-CNN and Prop.˜D.6 for SAE-FNO).
Fig. 2: SAE-CNN vs. SAE-FNO. a) Lifting accelerates learning. b) SAE-FNO’s superiority in recovering smooth concepts via truncated Fourier modes. c) Equivalent learning when SAE-FNO uses all Fourier modes and matched spatial receptive field of SAE-CNNs.
SAE-FNO We examined model recovery in function spaces. Our results show that: i) the lifting-induced preconditioning effect extends to L-SAE-FNO (Fig.˜2a, Prop.˜D.3); ii) SAE-FNO with truncated modes exhibits an inductive bias that favours recovery of smooth concepts, which is desirable when the underlying data concepts are smooth (Fig.˜2b); iii) SAE-FNOs can successfully infer representations from higher-resolution data, whereas SAEs fail to recover the true underlying representation when data resolution changes (Fig.˜3); and iv) when the integral operator is implemented as a convolution with weights spanning the full frequency modes and constrained to the same spatial support, the architectural inference and training dynamics of SAE-FNO (Prop.˜3.3) reduce to those of SAE-CNN, with weights related by a Fourier transform (Prop.˜3.4, Fig.˜2c).
Proposition 3.3 (Training Dynamics of SAE-FNO).
Consider an SAE-FNO whose integral operator is implemented as a convolution, with frequency-domain weights ${\bm{W}}_{c}$ defined as the Fourier transform of the SAE-CNN’s spatial kernels, i.e., ${\bm{W}}_{c}=\mathcal{F}{\bm{D}}_{c}$. The training dynamics of the SAE-FNO are then expressed as: ${\bm{D}}_{c}^{(k+1)}={\bm{D}}_{c}^{(k)}+\tfrac{\eta}{M}({\bm{x}}-\tfrac{1}{\sqrt{M}}\mathcal{F}^{-1}[{\textstyle\sum_{i=1}^{C}}\mathcal{F}{\bm{D}}_{i}^{(k)}\odot\mathcal{F}{\bm{z}}_{i,t}])\star{\bm{z}}_{c,t}$ where $M$ is the full number of modes, $\odot$ denotes element-wise multiplication, and $\star$ denotes correlation. If the SAE-FNO’s filters span the full frequency range and are constrained to have the same spatial support as in the SAE-CNN, then its dynamics reduce to those of the SAE-CNN (see Prop.˜3.4).
Proposition 3.4 (Training Dynamics of SAE-CNN and SAE-FNO).
The training dynamics of the SAE-FNO’s frequency-domain weights ${\bm{W}}_{c}^{(k+1)}={\bm{W}}_{c}^{(k)}+\tfrac{\eta}{\sqrt{M}}\mathcal{F}\left({\bm{x}}-\hat{{\bm{x}}}^{(k)}\right)\odot\mathcal{F}({\bm{z}}_{c,t})$ has an effective update in the original space, expressed as: ${\bm{D}}_{c}^{\prime(k+1)}={\bm{D}}_{c}^{\prime(k)}+\tfrac{\eta}{M}\left({\bm{x}}-\hat{{\bm{x}}}^{(k)}\right)\star{\bm{z}}_{c,t}$. If the architectures satisfy the inference equivalence condition (Prop.˜D.4, see  Prop.˜D.5 for lifting), then the dynamics are equivalent to that of a SAE-CNN: ${\bm{D}}_{c}^{(k+1)}={\bm{D}}_{c}^{(k)}+\eta({\bm{x}}-\sum_{i=1}^{C}{\bm{D}}_{i}^{(k)}*{\bm{z}}_{i,t})\star{\bm{z}}_{c,t}$.
(a) Original.
(b) $2\times$ Upsampling.
(c) $8\times$ Upsampling.
Fig. 3: SAE-FNO Upsampling Robustness Across Resolutions. SAE-FNO successfully infers the underlying sparse representations and reconstructs data across multiple discretization levels. The left panels show inference of 1-sparse code supports across 5 kernels, and the right panels display spatial-domain signal reconstruction (see also Fig.˜7).
Conclusion We introduced a unified framework for sparse model recovery to analyze representations across finite- and infinite-dimensional settings. Our analysis shows that the SAE-FNO’s inductive biases enable superior recovery of smooth concepts and generalization across resolutions where standard models fail. This provides a rigorous foundation for both model recovery in function spaces and mechanistic interpretability of large neural operators. While validated here on synthetic data, critical next steps include applying this framework to real-world settings, where based on the prior success of neural operators, we speculate that SAE-FNO will be highly effective in extracting meaningful concepts, including learning harmonic representations useful in scientific modelling and capturing global concepts or frequency-based coding in neuroscience.
Appendix A Appendix - Acknowledgments
A.S. conducted this work as a Dale and Suzanne Burger SURF Fellow through the Summer Undergraduate Research Fellowship (SURF) program at Caltech and gratefully acknowledges its funding. A.A. was supported by the Bren Chair Professorship and AI2050 Senior Fellowship.
B.T. proposed, designed, and led the project. A.S. implemented the methods and executed all experiments. A.S. and B.T. wrote the paper together. A.A. provided supervision, valuable feedback, and editorial comments.
Appendix B Appendix - Additional Background

### v3

@@ 3 Sparse Autoencoder Neural Operators
The goal of model recovery with SAE-NO in function spaces is to infer a sparse representation $\bm{z}$} as a sample from the function space $\mathcal{Z}$, representing the data sample from the function space $\mathcal{X}$ via $\mathcal{G}_{{\bm{D}}}$: $\text{(encoder)}\quad{\bm{z}}=\text{Enc}_{\theta}({\bm{x}}),\quad\text{(decoder)}\quad\hat{\bm{x}}=\text{Dec}_{{\bm{D}}}({\bm{z}})$, where $\text{Enc}_{\theta}$ is the encoder operator relating to $\mathcal{H}_{\theta}$, and model recovery involves estimating the concepts $\bm{D}$} parameterized within $\text{Dec}_{{\bm{D}}}$.
We consider a Fourier parameterization and define SAE Fourier Neural Operators (SAE-FNOs) as $\mathcal{G}_{{\bm{D}}}({\bm{z}})\coloneqq$ $\mathcal{F}^{-1}(\sum\nolimits_{c=1}^{C}{\bm{W}}_{c}\cdot(\mathcal{F}{\bm{z}}_{c}))$ where ${\bm{W}}_{c}=\mathcal{F}{\bm{D}}_{c}$ denotes the Fourier transform of the convolution kernel ${\bm{D}}_{c}$. SAE-NO is fundamentally different from all other SAEs to date; all SAEs parameterize the concepts in the same domain as the encoder input, mapping the representation via ${\bm{D}}{\bm{z}}$ to the reconstruction. SAE-NO, on the other hand, parameterize both encoding and decoding concepts as an integral kernel operator parameterized in Fourier space. We show that this Fourier and functional parameterization plays a critical role in shaping the characterization and generalization of the learned concepts. Owing to the duality between convolution in the spatial domain and multiplication in the Fourier domain, inference in SAE-FNOs can be viewed as recovering concepts under a sparse convolutional generative model (Definition A.1), implemented as a sum of convolutions between $\bm{z}$} and $\bm{D}$}. While SAE-FNO is general and can be applied to SAEs with different activation functions (e.g., ReLU), we introduce two distinct notions of sparsity for SAE-FNO: concept sparsity and domain sparsity.
(a) SAE-FNO.
(b) SAE-MLP.
(c) SAE-CNN.
(d) Concept correlation.
Figure 4: SAE-NO concept characterization and stability. We use the Hoyer score to characterize the locality and sparsity of the learned concepts, and examine the concept correlations (a) SAE-FNO shows stable concept characteristics across different concept sparsity levels. (b) In contrast, the characteristics of concepts learned by SAE change drastically as the sparsity level varies, indicating an undesirable sensitivity of concept properties to this choice. (c) Comparison of concept correlations for SAE-FNO and SAE-CNN: SAE-FNO exhibits a broader range of correlations, even in the presence of a shallow, orthogonality-inducing inductive bias.
(a) SAE-FNO (data size $8\times 8$).
(b) SAE-FNO (data size $32\times 32$).
(c) SAE-MLP (data size $8\times 8$).
(d) SAE-MLP (data size $32\times 32$).
Figure 5: Visualization of used concepts, across the full population, learned by SAE-FNO and SAE-MLP. From left to right, the concept sparsity is set to $25$, $50$, and $200$. (a–b) SAE-FNO exhibits stable concept characteristics across varying levels of concept sparsity, enabled by domain sparsity and the ability to adapt to different input domain sizes. As concept sparsity decreases, a larger number of concepts are activated across the dataset. For localized image patches ($8\times 8$), SAE-FNO learns sinusoidal patterns at varying frequencies, while for larger inputs ($32\times 32$), it learns localized frequency patterns such as edge detectors. (c–d) In contrast, the characteristics of concepts learned by SAE-MLP change drastically as concept sparsity increases, leading to a loss of interpretable structure. Moreover, SAE-MLP lacks the capacity to capture local spatial structure in the data. Hence, an important observation is that particularly, when data domain gets larger, the effective number of learned concepts in SAE-FNO is substantially smaller than in SAE-MLP, indicating the superior expressivity of SAE-FNO, particularly for inputs with large domains and localized structure.
Concept sparsity extends standard SAEs by enforcing that only a few concepts explain each data example, while domain sparsity captures the activation pattern within each concept’s feature map. As a result, concept sparsity induces group sparsity across feature maps, whereas domain sparsity enforces element-wise sparsity within each map. Intuitively, concept sparsity determines which concepts are active, while domain sparsity determines where they are expressed in the input domain. We show that this sparsity parameterization enables SAE-FNO to learn globally shared yet locally expressed concepts that are robust to changes in sparsity level. This enables SAE-FNO to represent datasets more compactly, using a smaller number of learned concepts.
For TopK as the activation function, SAE-FNO implements the following:
$\min_{{\bm{W}}\in\mathcal{W}}$ $\tfrac{1}{2}\|{\bm{x}}-\mathcal{F}^{-1}({\bm{W}}\cdot\mathcal{F}{\bm{z}}+{\bm{b}}_{\text{pre}})\|_{2}^{2}$ (3)
s.t $\bm{v}=(\mathcal{F}^{-1}(\bm{E}\cdot(\mathcal{F}{\bm{x}}-{\bm{b}}_{\text{pre}}))+{\bm{b}}_{\text{enc}})$
(concept sparsity) $\bm{a}=\text{TopK}\left([\|\bm{v}_{1}\|_{2},\cdots,\|\bm{v}_{C}\|_{2}]^{\top}\right)$
$\bm{v}=\mathbf{1}_{\bm{a}}\cdot\bm{v}$
(domain-sparsity) ${\bm{z}}_{i}=\text{TopK}\left(\bm{v}_{i}\right),\text{for}\ i\text{, where}\ \|\bm{v}_{i}\|_{2}>0$
We further extend this approach to concept recovery in lifted spaces (Definition 3.4), incorporating lifting and projection modules that are crucial components of neural operators.
Definition 3.4 (Lifted SAE-NO for Concept Recovery).
@@
The goal of model recovery with lifted SAE-NO (L-SAE-NO) in function spaces is to infer a sparse representation $\bm{z}$} as a sample from the function space $\mathcal{Z}$, representing the data from $\mathcal{X}$ via the operator $\text{Dec}_{{\bm{D}}_{{\bm{L}}}}$:
(lifting and encoder) ${\bm{z}}=\text{Enc}_{\theta}({\bm{y}}),\quad{\bm{y}}={\bm{L}}{\bm{x}}$ (4)
(decoder and projecting) $\hat{{\bm{x}}}={\bm{P}}\hat{{\bm{y}}},\quad\hat{{\bm{y}}}=\text{Dec}_{{\bm{D}}_{{\bm{L}}}}({\bm{z}}),$
@@
where $\text{Enc}_{\theta}$ encodes the lifted data $\bm{y}$}, and concept recovery is done via the learned lifted concepts ${\bm{D}}_{{\bm{L}}}$.
Proposition 3.5 shows that gradient dynamics of Fourier concepts $\bm{W}$} in L-SAE-FNO can be represented as concepts in the original and unlifted space as $\bm{D}$} where the lifting acts as a preconditioner, which if designed/learned properly can accelerate learning by inducing a more isotropic update. We show in Proposition C.8 that the dynamics of a L-SAE-FNO can be reduced to SAE-FNO when the lifting and projection are tied (${\bm{P}}={\bm{L}}^{\top}$) and the lifting operator is orthogonal (${\bm{L}}^{\top}{\bm{L}}={\bm{I}}$). This highlights that recovery of lifted concepts in relation to the original domain is possible. These results also hold for SAE-MLP and SAE-CNN (see Appendix).
Proposition 3.5 (Training Dynamics of Lifted SAE-FNO).
The training dynamics of the Lifted-SAE-FNO (L-SAE-FNO)’s frequency domain weights ${\bm{W}}_{L,c}^{(k+1)}={\bm{W}}_{L,c}^{(k)}+\dfrac{\eta_{L}}{\sqrt{M}}\mathcal{F}\left({\bm{L}}\left({\bm{x}}-\hat{{\bm{x}}}^{(k)}\right)\right)\odot\mathcal{F}({\bm{z}}_{c,k})$, with lifting operator $\bm{L}$} and projection $\bm{P}$}, has an effective update in the original space, expressed as: ${\bm{D}}_{c}^{(k+1)}={\bm{D}}_{c}^{(k)}+\dfrac{\eta_{L}}{M}\left(({\bm{L}}^{\top}{\bm{L}})\left({\bm{x}}-\hat{{\bm{x}}}^{(k)}\right)\right)\star{\bm{z}}_{c,t}$, where ${\bm{L}}^{\top}{\bm{L}}$ acts as a preconditioner, potentially. If the architectures satisfy the architectural inference equivalence condition (Proposition C.8), then the dynamics is equivalent to that of an SAE-FNO: ${\bm{D}}_{c}^{\prime(k+1)}={\bm{D}}_{c}^{\prime(k)}+\dfrac{\eta}{M}\left({\bm{x}}-\dfrac{1}{\sqrt{M}}\mathcal{F}^{-1}\left(\sum\limits_{i=1}^{C}\mathcal{F}({\bm{D}}_{i}^{(k)})\odot\mathcal{F}({\bm{z}}_{i,k})\right)\right)\star{\bm{z}}_{c,k}$.
(a) SAE-FNO ($8\times 8)$.
(b) SAE-MLP ($8\times 8$).
(c) $\text{R}^{2}$ ($8\times 8$).
(d) SAE-FNO ($32\times 32)$.
(e) SAE-MLP ($32\times 32$).
(f) $\text{R}^{2}$ ($32\times 32$).
Figure 6: Concept utilization and expressivity by SAE-FNO and SAE-MLP. (a,b,d,e) The $\ell_{0}$ code statistics (y-axis) across the data population are plotted against concept indices, sorted by decreasing $\ell_{0}$ utilization (x-axis). (c,f) Goodness-of-fit ($\text{R}^{2}$). SAE-FNO shows a much sharper decrease in concept utilization as a function of number of concepts compared to SAE-MLP. The area under the curve (concept utilization) quantifies the number of learned concepts required to characterize the data. The flatter curve for SAE-MLP suggests that as data domain grows, SAE-MLPs require a much larger number of concepts to achieve the same representation capability as SAE-FNO. This implies that the concepts learned by SAE-FNO are more generalizable and used more globally and efficiently across the dataset. The $R^{2}$ results demonstrate the limited capacity of SAE-MLP to capture the underlying local sturcture of the data population.
@@ 4 Results
In model interpretability, SAEs are commonly used in two settings. One approach trains SAEs on internal model representations to identify latent concepts (Huben et al., 2023), but this requires additional post-processing to map internal features to human-interpretable attributes, introducing practical complexity. Alternatively, recent work trains SAEs directly on samples from the model’s output distribution, operating on human-interpretable signals rather than internal activations (Bohacek et al., 2025; Jiang et al., 2025). This paradigm is particularly attractive for generative models, as it enables direct analysis of generated data and learned concepts without explicit alignment steps. We adopt this second approach and train SAE-FNOs directly on data with local, temporal, and spatial structure, comparing them to standard SAEs to enable immediate interpretation of learned concepts in the data domain.
Figure 7: Generalization of SAE concepts under data distribution shift. Under a distribution shift induced by data rotation, the reconstruction error of SAE-MLP increases sharply relative to in-distribution performance. In contrast, SAE-FNO learns more generalizable concepts, leading to a substantially smaller increase in reconstruction error on out-of-distribution samples. This behavior is consistent across a wide range of Fourier modes ($15$, 2020, and $28$) used in the architecture.
@@
Concept Recovery Under Known Generative Model. We first characterize SAE-FNO under a known ground-truth generative model. We instantiate the model with an unrolled encoder (Tolooshams and Ba, 2022) (see Appendix for details). As shown in Figure 2(a), SAE-FNO is successful in identifying the underlying generative model and learning the ground-truth concepts. We further analyze concept recovery in the lifted space, and found: i) underlying concepts can be recovered in the lifted space; ii) lifting can act as a preconditioner (Proposition 3.5), accelerating recovery (Figure 2(b)) by reducing dictionary atom correlation, promoting a more isotropic loss landscape (Figure 10); ; iii) when lifting and projection are tied and orthogonal, the architectural inference (Proposition C.8) and training dynamics (Proposition 3.5) of L-SAE-NO reduces to SAE-NO.
SAE-NO Discretization Capability. Conventional SAEs operate on fixed-grid data in Euclidean spaces such as $\mathbb{R}^{n}$, with representations tied to a specific resolution. In contrast, the functional parameterization of SAE-FNOs inherits key advantages of neural operators, enabling analysis and inference across discretization levels. We demonstrate this capability under a known generative model (Figure 3), where SAE-FNO, unlike SAE-CNN, generalizes to data at higher resolutions than those seen during training. We further validate this property on MNIST (Figure 11): an SAE-NO with $p=1000$ concepts trained at $28\times 28$ continues to represent data accurately up to $112\times 112$, while the reconstruction error of a standard SAE increases sharply beyond the training resolution. This discretization generalization persists across different frequency truncations.
SAE-NO Concept Stability and Correlation A defining characteristic of SAE-NOs that distinguishes them from existing SAEs lies in their concept parameterization. In particular, SAE-FNOs parameterize concepts in the Fourier domain. We characterize the learned concepts using the Hoyer score, a normalized measure for sparsity in $[0,1]$, where 1 indicates a fully sparse concept and 0 indicates a fully dense concept. We compute the score for each concept individually. For a flattened representation $v_{c}\in\mathbb{R}^{N}$ of the $c$-th concept, the Hoyer score is given by:
$\text{Hoyer}(v_{c})=\frac{\sqrt{N}-\frac{\|v_{c}\|_{1}}{\|v_{c}\|_{2}}}{\sqrt{N}-1}$ (5)
We report the distribution of these scores over all $p$ concepts to characterize the sparsity of the learned dictionary.
As shown in Figure 4a, this parameterization offers concepts whose characteristics, quantified by the Hoyer score, remain stable and robust across different choices of sparsity level in the representation. In contrast, the characteristics of concepts learned by standard SAEs (Figure 4b) vary substantially as the sparsity level changes. We emphasize that such stability is a crucial and desirable property in practical applications: when learned concepts are highly sensitive to the choice of sparsity level, it becomes unclear which concepts can be reliably interpreted or trusted by the user.
Another important property of SAE-NOs is their flexibility in learning correlated concepts. This contrasts with the inductive bias of shallow SAEs, which tend to learn approximately orthogonal concepts (Costa et al., 2025). While recent work has begun to move beyond strict orthogonality, it remains desirable to develop SAEs that can learn non-orthogonal concepts even in the presence of orthogonality-inducing biases. Figure 4d illustrates this capability for SAE-FNOs compared to CNN-based counterparts. The histogram of pairwise concept correlations for SAE-FNOs exhibits a broader distribution, whereas SAE-CNNs show a sharply peaked distribution centered near zero. This increased flexibility stems from the functional parameterization of SAE-FNOs, which associates each concept with a feature-map representation. As a result, SAE-FNOs can learn concepts with similar overall code energy that differ in how and where they are expressed across the input domain, for example by appearing in different local regions.
Figure 8: Lifting accelerates and improves performance. Lifting acts as a preconditioner introducing an implicit acceleration to learning. In real-world setting, this can result in better fit of data by reducing the impact of noise on learned concepts. The impact of lifting also extends to conventional SAEs (left) SAE-FNO vs. Lifted SAE-FNO. (right) SAE vs. Lifted SAE.
Concept Visualization and Utilization. We visualize the learned concepts on natural images and examine their utilization across the data population (see Appendix for detailed experimental setup). We particularly show how conventional SAEs fail when the underlying model/data exhibit spatially local structure, which cannot be efficiently extracted by SAE-MLPs. Figure 5 shows that the SAE-FNO learns sinusoidal filters for local patches and localized frequency patterns for larger images. While such concept characteristic of SAE-FNO is stable against the concept sparsity level, concepts learned by SAE-MLP change drastically as concept sparsity increases, losing interpretable structure.
This qualitative difference directly explains the efficiency observed in Figure 6, where we examine concept utilization across the population. SAE-FNO exhibits a significantly sharper decrease in concept usage across the data population, implying that a smaller, compact set of these structured concepts is sufficient to characterize the dataset. In contrast, SAE-MLP shows a much flatter distribution, requiring a significantly larger number of active concepts. While both models may achieve comparable reconstruction ($R^{2}$) on small domains (patches), SAE-FNO outperforms SAE-MLP on larger image domain, due to the need to extract local structures for generalization. Additionally, we argue that the primary advantage of SAE-FNO lies in the characteristics of the learned concepts and their interpretability. Consequently, our analysis focuses on these properties, such as stability, resolution invariance, and structure, and we prioritize these over reconstruction metrics.
Lifting Accelerated Learning and Improved Performance. An additional contribution of our work is the ability to lift an SAE into a latent space and perform concept and sparse representation learning in this lifted domain. Complementing our earlier lifted results, we now shows that, beyond faster convergence, lifting provides two key benefits (Figure 8): (i) the acceleration effect extends beyond operator-based SAEs and applies to non-operator SAEs, and (ii) lifting improves concept learning, leading to lower reconstruction error. This gain can be interpreted through mechanisms akin to the implicit regularization of gradient descent in deep models (Neyshabur, 2017; Arora et al., 2019). From a concept learning perspective, lifting shifts stochasticity and noise to the lifted space, while linear projection back to the original domain induces a smoothing effect that facilitates the extraction of more robust, higher signal-to-noise concepts.
Concept Generalization Under Distribution Shift. Finally, we show that SAE-FNO learns more generalizable concepts than SAE-MLP under a distribution shift induced by image rotation (Figure 7). While standard SAEs exhibit a substantial increase in reconstruction error, SAE-NOs show a much smaller relative increase in mean squared error, indicating more stable and representative concepts. We note that this experiment provides preliminary insight into concept generalization under distribution shift; a comprehensive ablation study of out-of-distribution generalization is left for future work.
5 Conclusion
We introduced sparse autoencoder neural operators (SAE-NOs), a framework generalizing sparse representation learning to infinite-dimensional function spaces. By parameterizing concepts in the Fourier domain via SAE-FNO, we move beyond scalar activations to feature-map representations that explicitly encode where and how concepts are expressed through domain sparsity. We demonstrated that this functional parameterization fundamentally enhances concept learning: SAE-NOs exhibit resolution invariance, robust recovery under distribution shifts, and stability in learned concept characteristics across varying sparsity levels. Our findings highlight parameterization as crucial in shaping learned representations. Moreover, we showed that the lifting mechanism acts as a preconditioner, accelerating recovery and improving robustness to noise, enabling high-SNR concept learning.
In this work, we focus on vision datasets with local structure and perform concept learning in a controlled setting to better characterize this new model. While our primary goal is to analyze the concepts learned by SAE-NOs and demonstrate that functional parameterization fundamentally improves concept learning in sparse autoencoders, several limitations remain. In particular, our empirical evaluation is restricted to controlled vision experiments and moderate-scale real data; extending SAE-NOs to large-scale, high-dimensional settings, such as foundation model activations or complex scientific simulations, requires further investigation.
Impact Statement
This work advances the field of machine learning by introducing a new framework for representation learning and interpretability based on sparse autoencoder neural operators (SAE-NO). By extending sparse autoencoders to functional and multi-resolution settings, our approach enables more stable, generalizable, and interpretable concept representations, with potential benefits for scientific modelling and mechanistic interpretability in general.
The methods proposed here are intended for understanding and characterizing learned representations rather than for deployment in high-stakes decision-making systems. hence, we do not anticipate immediate negative societal impacts. However, as with any interpretability or representation learning tool, insights derived from these models could influence how complex systems are analyzed or trusted. Care should therefore be taken when applying SAEs in sensitive domains, and conclusions should be interpreted in conjunction with domain expertise.
Overall, this work aims to contribute foundational insights into representation learning and interpretability. We do not identify any foreseeable ethical concerns specific to this contribution.
Appendix A Appendix - Additional Background